### v2

# Gauge Flow Models

Abstract
This paper introduces Gauge Flow Models, a novel class of Generative Flow Models. These models incorporate a learnable Gauge Field within the Flow Ordinary Differential Equation (ODE). A comprehensive mathematical framework for these models, detailing their construction and properties, is provided. Experiments using Flow Matching on Gaussian Mixture Models demonstrate that Gauge Flow Models yield significantly better performance than traditional Flow Models of comparable or even larger size. Additionally, unpublished research indicates a potential for enhanced performance across a broader range of generative tasks.
1  Introduction
A Gauge Flow Model is defined within the mathematical framework of associated bundles:
$\hat{A}=P\times_{G}F$
where $P=(P,M,G,\pi_{P})$ is a principal bundle with structure group $G$ and $F$ is a $G$-space. The dynamics are governed by the following neural ODE:
$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))$
This equation incorporates several novel key components:
  • $v_{\theta}(x(t),t)\in TM$ is a learnable vector field, defined as a section of the tangent bundle over the base manifold $M$.
  • $\alpha(t)$ represents a learnable schedule.
  • $A_{\mu}(x(t),t)$ is a learnable Gauge Field valued in the Lie algebra of the Gauge Group $G$.
  • $d^{\mu}(x(t),t)$ is a direction vector field, modeled as a smooth section of the tangent bundle $TM$.
  • $v_{s}(x(t),t)$ is a further learnable vector field corresponding to a section on the associated bundle $P\times_{G}F$.
  • $\Pi_{M}:P\times F\to TM$ is a smooth, learnable projection map.
The novelty of this Flow ODE lies in the inclusion of the Gauge Term, which introduces a novel combination of learnable and non-learnable vector fields – a feature absent from standard Neural Flow Models. The standard Neural Flow Model, as defined by [1, 2], is characterized by:
$\frac{dx(t)}{dt}=v_{\theta}(x(t),t)$
The Gauge Field associated with a predefined Gauge Group (e.g., $SO(N)$ or $SU(N)$) establishes a crucial geometric inductive bias. By explicitly incorporating this bias, the model learns to represent data more efficiently, favoring arrangements that align with the imposed symmetries. Consequently, these models often achieve stronger performance and improved robustness compared to models lacking this geometric constraint. This approach is particularly advantageous in domains like protein or drug design, where molecules frequently exhibit rotational or translational symmetries.
2  Mathematical Background
This section introduces the mathematical foundations of Gauge Flow Models, primarily utilizing the concept of Fiber Bundles. First, the necessary mathematical tools are established, culminating in a comprehensive description of Gauge Flow Models. A working familiarity with concepts from Differential Geometry – including manifolds, vector fields, differential forms, and Lie groups – is assumed. For readers seeking a refresher or introduction to these topics, several relevant resources are listed below.
  • Differential Geometry: [3, 4, 5, 6, 7]
  • Gauge Theory: [8, 9, 10]
  • Lie Groups and Lie Algebras: [11, 12]
2.1  Fiber Bundle
A fiber bundle is a geometric construction that assembles a continuous family of fibers, one for each point of a base space, in a way that varies continuously across the base. Locally, it resembles a product of the base space and the fiber. However, it’s important to note that fiber bundles can generally deviate from this local product space in non-trivial ways. An illustration of a fiber bundle is provided below.
Figure 1: A fiber bundle can be visualized as a base space with a fiber attached to each point. As shown, the base space corresponds to the plane and the blue lines represent the fibers.
A fiber bundle is formally defined as the data $(E,B,F,\pi)$, where:
  • Total space: E is the union of all the fibers.
  • Base space: B is the parameter space.
  • Fiber: F is the typical space sitting over each point of B.
  • Projection: $\pi:E\mapsto B$ is a continuous surjective map.
Ensuring local triviality – meaning that the bundle behaves like a product locally – requires that for every point $x\in B$ there exists an open neighborhood $U_{x}\subset B$ and a fiber preserving homeomorphism $\Psi$:
$\Psi:\pi^{-1}(U_{x})\mapsto U_{x}\times F$
such that the following diagram commutes:
$\pi^{-1}(U_{x})$$U\times F$$U_{x}$$U_{x}$$\scriptstyle{\Psi}$$\scriptstyle{\pi}$$\scriptstyle{\mathrm{pr}_{1}}$$\scriptstyle{\mathrm{id}_{U_{x}}}$
This local triviality condition ensures that the bundle can be locally described as a product of the base space and the fiber.
2.1.1  Connection on a Fiber Bundle
A connection on a fiber bundle defines a rule for parallel transport, enabling the comparison of fibers at infinitesimally nearby points. This provides a way to transport elements along curves in the base manifold in a geometrically consistent manner.
Figure 2: A connection on a fiber bundle defines a mapping that relates each fiber to its infinitesimally near neighboring fibers. This enables local ‘parallel transport’ within the bundle.
Mathematically, a connection on a fiber bundle $(E,B,F,\pi)$ is defined as a smoothly varying family of horizontal subspaces $H_{p}E$:
$HE\subset TE$
where:
  • 1.)
    At every point $p\in E$, the tangent space $T_{p}E$ is split into a horizontal component $H_{p}E$ and a vertical component $V_{p}E$:
    $T_{p}E=H_{p}E\oplus V_{p}E$
    The vertical component, $V_{p}E$ is defined via the differential $d\pi$:
    $V_{p}E:=ker(d\pi_{p}:T_{p}E\mapsto T_{\pi(p)}B)$
  • 2.)
    The differential of the projection map $d\pi_{p}$, restricted to the horizontal subspace $H_{p}E$ is a linear isomorphism:
    $d\pi_{p}|_{H_{p}E}:H_{p}E\rightarrow T_{\pi(p)}B$
In essence, a connection is precisely the choice of such a smoothly varying family of horizontal subspaces $H_{p}E\subset T_{p}E$. This particular type of connection is often referred to as an Ehresmann Connection.
2.1.2  Principal Bundle
[…]
$\bullet:$ $G\times G\rightarrow G$
$(g,h)\mapsto g\bullet h$
such that the following group axioms are satisfied:
  • Closure:
    $g\bullet h\in G\quad\forall\,g,h\in G$
  • Associativity:
    $g\bullet(h\bullet f)=(g\bullet h)\bullet f\quad\forall\,g,h,f\in G$
  • Identity Element:
    $g\bullet e=g=e\bullet g\quad\exists\,e\in G,\;\forall g\in G$
  • Inverse Element:
    $g\bullet g^{-1}=e=g^{-1}\bullet g\quad\exists g^{-1}\in G,\;\ \forall g\in G$
Moreover, the group operation $\bullet$ must satisfy the smoothness condition:
$G\times G\rightarrow G$
$(g,h)\mapsto g\bullet h^{-1}$
[…]
Lie Algebra A Lie algebra $\mathfrak{g}$ of a Lie Group $G$ is simply the tangent space $T_{e}G$ at the identity element $e$ to the Lie Group. This vector space $\mathfrak{g}$ inherits a bilinear bracket [,][-,-] from the group structure of $G$:
$\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$
that satisfies the following conditions for all $X,Y,Z\in\mathfrak{g}$:
  • Antisymmetry:
    [X,Y]=[Y,X][X,Y]=-[Y,X]
  • Jacobi identity:
    $[X,[Y,Z]]+[Y,[Z,X]]+[Z,[X,Y]]=0$
These two structures – the Lie algebra and the Lie Group – allow the definition of a principal bundle and, subsequently, a connection on it.
A principal bundle $(P,B,G,\pi)$ is a smooth fiber bundle characterized by:
  • Fiber: The typical fiber is the Lie group $G$.
  • Right Action: A smooth right action of $G$ on $P$: $P\times G\rightarrow P$ defined by $p\cdot g$ for all $p\in P$ and $g\in G$.
This action is subject to the following constraints:
  • Fiber-preserving:
    $\pi(p\cdot g)=\pi(p)\quad\forall\,p\in P,\;\forall g\in G$
  • Free:
    $p\cdot g=p\;\Longrightarrow\;g=e\quad\forall\,p\in P,\;g\in G$
  • Transitive on each fiber:
    $\forall\,b\in B,\;\forall\,p,q\in\pi^{-1}(b)\;\exists!\,g\in G\text{ such that }p\cdot g=q.$
[…]
Connection on a Principal Bundle A connection on a principal bundle $P=(P,B,G,\pi)$ is a connection $HP\subset TP$ on the fiber bundle $P$ characterized by the G-equivariance condition:
$dR_{g}\!\bigl{(}H_{p}P\bigr{)}\;=\;H_{\,p\cdot g}P\qquad\forall\,p\in P,\;\forall\,g\in G,$
where $R_{g}$ is the right translation defined as:
$R_{g}:P\rightarrow P,\;R_{g}(p):=p\cdot g$
with its differential, $dR_{g}$, is the pushforward:
$dR_{g}|_{p}:T_{p}P\rightarrow T_{p\cdot g}P$
This can equivalently be expressed in terms of a connection 1-form $\omega$.
$\omega\in\Omega^{1}(P)\otimes\mathfrak{g}$
where the kernel at each point is exactly the horizontal subspace:
$H_{p}P=ker(\omega_{p})$
This $\omega$ must satisfy the following two compatibility conditions:
  • Reproduction on verticals: For each $X\in\mathfrak{g}$, the fundamental (vertical) vector field $X^{\#}$ (the infinitesimal generator of the G-action) obeys:
    $\omega(X^{\#})=X$
  • Ad-equivariance:
    $R^{\ast}_{g}\omega=Ad(g^{-1})\omega$
    where $R^{\ast}_{g}$ is the pullback of $R_{g}$ and $Ad(g^{-1})$ is the adjoint action of the Lie group $G$ on the Lie algebra $\mathfrak{g}$.
The Gauge Field $A$ can be defined by selecting a local section $s$ (a gauge choice) on an open neighborhood $U$:
$s:U\subset B\rightarrow P$
where the section is defined such that its projection onto $U\subset M$ recovers the identity on $U$:
$\pi_{P}\circ s=Id_{U}$
In this case, the Gauge Field is given as the pullback of the connection 1-form $\omega$:
$A:=s^{\ast}\omega\in\Omega^{1}(U)\otimes\mathfrak{g}$
Given a $G$-valued function $g:U\rightarrow G$, the local section is transformed as:
$s^{\prime}(x)=s(x)g(x)\quad\forall x\in U$
This transformation induces a gauge transformation on the gauge field $A$ given by:
$A^{\prime}(x)=g^{-1}(x)A(x)g(x)+g^{-1}(x)dg(x)$
where:
$A=s^{\ast}\omega\quad A^{\prime}=s^{\prime\ast}\omega$
The field strength (or curvature) $F\in\Omega^{2}(U)\otimes\mathfrak{g}$ of a gauge field $A$ is defined as:
$F=dA+A\wedge A$
For further information on differential forms and pullbacks, consult the references listed above.
2.1.3  Associated Bundle
An associated bundle $\hat{A}$ is constructed by “gluing” together copies of a smooth manifold $F$ using the transition data of a principal bundle $P$, in such a way that each fiber inherits a given $G$-action. Importantly, any connection on the principal bundle induces a natural notion of parallel transport on the associated bundle, thereby defining a connection on $\hat{A}$ as well. This allows for a meaningful comparison of nearby fibers.
Mathematically, the associated bundle $\hat{A}$ is defined as the quotient:
$\hat{A}=P\times_{G}F=(P\times F)/\sim$
with principal bundle $P=(P,B,G,\pi)$ and $F$ being a smooth manifold (or vector space) equipped with a smooth left action of Lie group $G$.
The associated bundle $\hat{A}$ is specified by the following:
  • Diagonal $G$-action: The (right) $G$-action on $P\times F$ defined as:
    $(p,f)\cdot g\;=\;\bigl{(}p\cdot g,\;g^{-1}\!\cdot\!f\bigr{)}$
    This action is free and proper.
  • Equivalence relation: The equivalence relation $\sim$ is defined as:
    $(p,\;f)\sim(p\!\cdot\!g,\;g^{-1}\!\cdot\!f)\qquad\forall\,p\in P,\;g\in G,\;f\in F.$
  • Quotient Map: The quotient map $q$ is specified by:
    $q:P\times F\rightarrow\hat{A},\;(p,f)\mapsto[p,f]$
  • Local trivialisations: For every open set $U\subset B$ with a local section $s:U\to P$, the map
    $\Phi_{s}:U\times F\longrightarrow\pi_{\hat{A}}^{-1}(U),\quad(x,f)\longmapsto[s(x),f]$
    is a diffeomorphism. The associated bundle projection $\pi_{\hat{A}}$ is defined as:
    $\pi_{\hat{A}}:\hat{A}\rightarrow B,\;\pi_{\hat{A}}([p,f])=\pi(p)$
    Consequently, $\hat{A}$ is locally isomorphic to the product bundle $U\times F$. When $F$ is a vector space and the $G$-action is linear, these trivialisations endow $\hat{A}$ with the natural structure of a vector bundle.
The construction of an associated bundle can be illustrated by the following commutative diagram:
[…]
where $pr_{P}:P\times F\rightarrow P$ represent the projection onto the first factor.
3  Gauge Flow Model
A Gauge Flow Model is defined within the framework of the associated bundle $\hat{A}$:
$\hat{A}=P\times_{G}F$
with the following:
  • The principal bundle $P=(P,M,G,\pi_{P})$ with Lie group $G$ and base manifold $M$.
  • The typical fiber space $F$ with a smooth left action of the Lie group $G$.
The dynamics of a Gauge Flow Model are governed by the subsequent differential equation:
$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))$
Here, the components are defined as follows:
  • $v_{\theta}(x(t),t)\in TM$ is a learnable vector field modeled by a Neural Network.
  • $\alpha(t)$ is a time dependent weight, also modeled by a Neural Network.
  • AA is the learnable gauge field corresponding to the connection on the principal bundle $P$, also represented by a Neural Network.
  • $d^{\mu}(x(t),t)\in TM$ is the direction vector field.
  • $v_{s}(x(t),t):M\rightarrow P\times_{G}F$ is the fiber section, which can be learned by a Neural Network.
  • $\Pi_{M}:P\times F\rightarrow TM$ is a smooth projection from the product bundle $P\times F$ to the tangent bundle $TM$ of the base manifold $M$. This projection is covered, ensuring that the projection of any point $(p,v)$ from the product bundle collapses to the same base point $p$ within $M$:
    $\pi_{TM}(\Pi_{M}(p,v))=\pi_{P}(p)\quad\forall(p,v)\in P\times F$
The Gauge Flow Model is defined for any differentiable base manifold $M$. However, training the model requires $M$ to admit the structure of a Riemannian manifold, specifically possessing a Riemannian metric $g$ defined on it.
Training a Gauge Flow Model employs the Riemannian Flow Matching (RFM) framework [13], which generalizes ordinary Flow Matching (FM) [2]. The training objective is the Gauge Flow Model loss, defined as:
$\mathcal{L}_{\mathrm{GFM}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x\sim p_{t}\end{subarray}}\bigg{\lVert}\Big{[}v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))\Big{]}-u_{t}(x)\bigg{\rVert}_{g_{x}}^{2}$
where $\|\cdot\|_{g_{x}}$ is the norm induced by the Riemannian metric $g$ of the base manifold $M$ at point $x$. Here, both terms lie in the tangent space of the base manifold $M$:
$\Big{[}v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))\Big{]}\in T_{x}M$
$u_{t}(x)\in T_{x}M$
The Gauge Flow Model loss incorporates the following components:
  • Probability-density path $p_{t}:M\to\mathbb{R}_{+}$, $t\in[0,1]$, as defined in [13]. For each $t$, $p_{t}$ satisfies
    $\int_{M}p_{t}(x)\,d\mathrm{vol}_{x}=1,$
    where $d\mathrm{vol}_{x}$ is the Riemannian volume form.
  • Target vector field $u_{t}(x)$, obtained from the conditional field $u_{t}(x\mid x_{1})$ by
    $u_{t}(x)=\int_{M}u_{t}\!\bigl{(}x\mid x_{1}\bigr{)}\,\frac{p_{t}\!\bigl{(}x\mid x_{1}\bigr{)}\,q(x_{1})}{p_{t}(x)}\,d\mathrm{vol}_{x_{1}},$
    with $q(x)\coloneqq p_{t=1}(x)$ and $x_{1}$ denoting the sample at time $t=1$.
Computing the Gauge Flow Model (GFM) loss $\mathcal{L}_{\mathrm{GFM}}$ exactly is generally intractable. However, this loss is equivalent to the Riemannian Flow Matching (RFM) objective. To address this intractability, [13] introduces the Riemannian Conditional Flow Matching (RCFM) loss—a single-sample, unbiased Monte Carlo estimator of the RFM objective. The key idea is to replace the marginal target field $u_{t}(x)$ with its conditional counterpart $u_{t}(x\mid x_{1})$, enabling tractable training. This formulation allows for efficient, simulation-free learning on simple geometries and scalable neural network training on general Riemannian manifolds. Notably, this approach can also be applied to train Gauge Flow Models.
For full derivations and implementation details, refer to [13, 2].
4  Experiments
This study compares the novel Gauge Flow Models to standard Flow Models using a generated Gaussian Mixture Model (GMM) dataset. The Gauge Flow Models were evaluated for several values of $N$ associated with the Lie Group $G=SO(N)$.
4.1  Model
The configuration of the Gauge Flow Models consists of the following components:
  • The trivial principal bundle $P=(P=M\times G,M,G,\pi)$ with structure Lie group $G=SO(N)$ and base manifold $M=\mathbb{R}^{N}$.
  • The typical fiber $F=\mathbb{R}^{N}$.
  • The vector field $v_{\theta}(x(t),t)$ defined on the tangent bundle $TM$, is modeled by a Neural Network.
  • The alpha function $\alpha(t)$ is also modeled using a Neural Network.
  • The gauge field $A\in\Omega^{1}(M)\otimes\mathfrak{g}$, valued in the Lie algebra $\mathfrak{g}=\mathfrak{so}(N)$ of the Lie group $G=SO(N)$, is represented by a Neural Network.
  • The directional vector field $d^{\mu}(x(t),t)$ is either identified with the vector field $v_{\theta}(x(t),t)$ or $v_{s}(x(t),t)$ leading to two distinct Gauge Flow Model variants.
  • The vector field $v_{s}(x(t),t):M\rightarrow P\times_{G}F$ is expressed in the gauge defined by orthonormal bases of $\mathbb{R}^{N}$. In this gauge, the field $v_{s}(x(t),t)$ corresponds to a global map $M\rightarrow F$, which is modeled using a Neural Network.
  • The mapping $\Pi_{M}:P\times\mathbb{R}^{N}\rightarrow M\times\mathbb{R}^{N}$ acts as the identity function on both $M$ and $\mathbb{R}^{N}$, with the $G$-component effectively projected out as a consequence of the chosen gauge.
The Neural Network models used are standard Multi-Layer Perceptrons (MLPs) with the subsequent configurations:
Field Layer Dimensions Activation Function
$v_{\theta}(x(t),t)$ [N + 1, 32/64, 32/64, N] SiLU
AA [N + 1, 32/64, 32/64, $N^{2}(N-1)//2$ ] SiLU
$v_{s}(x(t),t)$ [N + 1, 32/64, 32/64, N] SiLU
$\alpha(t)$ [1, 16, 1 ] SiLU
The layer width depends on the dimension $N$ and is set to 32 for $N>10$ and to 64 otherwise. Here, SiLU refers to the Sigmoid Linear Unit activation function. The Plain Flow Model is implemented as a multilayer perceptron (MLP), where the number of layers and their widths are functions of the input dimension $N$. Initially, the network architecture follows the layer dimensions $[N+1,128,128,128,N]$, with SiLU used as the activation function throughout. These Gauge Flow Models are generally trained within the framework of Riemannian Flow Matching [13]. However, since the base manifold is $\mathbb{R}^{N}$, training can be performed using the standard Flow Matching [2].
4.2  Dataset
The datasets used in the experiment are generated Gaussian-Mixture Model (GMM) datasets with:
$n_{\text{train}}=15{,}000,\qquad n_{\text{test}}=5{,}000$
i.i.d. samples from a mixture of $K=3000$ equally-weighted Gaussian components in ambient dimension $N\in\{3,\dots,32\}$. The GMM datasets have the following specifications:
  • Mixture weights: $\pi_{k}=\frac{1}{K},\;k=0,\dots,K-1$.
  • Covariances: $\Sigma_{k}=0.5\,I_{N}$ (isotropic).
  • Spread parameter: $\alpha=25$.
  • Component means: $\mu_{k}$ for each component index $k$
    1. 1.
      Primary axis: $a_{1}=k\bmod N,\;\mu_{k,a_{1}}=(-1)^{k}\,\alpha.$
    2. 2.
      Secondary axis: $a_{2}=(k+\lfloor K/2\rfloor)\bmod N.$ If $a_{2}\neq a_{1}$ set $\mu_{k,a_{2}}=(-1)^{k+1}\,\tfrac{1}{2}\alpha.$
    3. 3.
      Extra offset (only when $K>N$ and $k\geq N$):
      let $b=(a_{1}+\lfloor k/N\rfloor)\bmod N$ and add
      $\mu_{k,b}\;{+}=\;s_{k}\,0.1\,\alpha\,\lfloor k/N\rfloor,\qquad s_{k}=\begin{cases}+1,&k\bmod 3=0,\\ -1,&\text{otherwise}.\end{cases}$
Sampling:  Draw $k\sim\mathrm{Cat}(\pi)$ and $x\sim\mathcal{N}\!\bigl{(}\mu_{k},\,0.5\,I_{N}\bigr{)}.$
4.3  Results
The performance of the models on the GMM dataset was evaluated using both training and testing loss, along with the number of parameters.
4.3.1  Train Loss
The Gauge Flow Model utilizing the directional vector field $v_{\theta}(x(t),t)$ or $v_{s}(x(t),t)$, demonstrated superior training performance. To facilitate meaningful comparisons, the training loss was normalized relative to the loss of the $v_{\theta}(x(t),t)$ Gauge Flow Models. Notably, the training performance varied across different Lie group dimensions ($N$s), with the Gauge Flow Model using $v_{\theta}(x(t),t)$ consistently exhibiting the best results for most values of $N$.
Figure 3: Train Loss Comparison (lower is better): The normalized train loss is plotted against the dimension NN to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v_{s}(x(t),t)$ with the standard plain Flow Model.
4.3.2  Test Loss
The test loss exhibited similar performance trends to the train loss. Both Gauge Flow Model variants significantly outperformed the standard Flow Model across all values of $N$.
Figure 4: Test Loss Comparison (lower is better): The normalized test loss is plotted against the dimension NN to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v_{s}(x(t),t)$ with the standard plain Flow Model.
4.3.3  Number of Parameters
As shown in Figure 5, the plain Flow Model had slightly more parameters than the Gauge Flow Models across all dimensions $N$, but the two variants ultimately had the same number of parameters.
Figure 5: Number of Parameters: A comparison of the number of parameters for the plain Flow Model and the Gauge Flow Models across various dimensions NN.
5  Related Work
The concept of Continuous Normalizing Flows (CNFs), as introduced in [1], traditionally relies on computationally intensive Ordinary Differential Equation (ODE) solvers during training, guided by a maximum likelihood objective. While effective, this training procedure scales poorly due to the repeated integration of ODEs. To mitigate this limitation, the Flow Matching paradigm was proposed in [2]. Instead of maximizing likelihood, Flow Matching formulates training as an $L_{2}$ regression problem that directly aligns a learnable vector field with a target velocity field. This target velocity is derived from known boundary conditions, significantly reducing computational overhead during training. Initially formulated in Euclidean space, Flow Matching was later extended to Riemannian manifolds in the Riemannian Flow Matching (RFM) framework [13]. In RFM, the vector field matching operates on curved spaces while respecting the geometry induced by a fixed Riemannian metric. However, it is important to note that the connections—i.e., the geometric structures defining parallel transport and differentiation—are determined by the manifold and are not themselves learnable parameters.
Parallel to these developments, several lines of research have explored the connections between gauge theory and deep learning. A notable example is presented in [14], where a gauge-invariant framework is introduced that encodes invariance directly into the inputs by constructing gauge-invariant features before passing them through standard neural architectures. This approach imposes gauge symmetry directly on the inputs instead of embedding it in the model dynamics.
Another approach leveraging geometric and gauge-theoretic principles is presented in [15], where standard Convolutional Neural Networks (CNNs) are generalized to operate on arbitrary manifolds by enforcing coordinate independence. This work builds upon the foundations laid by steerable CNNs [16] and Group Equivariant CNNs (G-CNNs) [17], extending the principle of equivariance to manifold-valued data. Further advancing this direction, Lattice Gauge Equivariant Convolutional Neural Networks (L-CNNs) [18] incorporate local gauge symmetry directly into the network architecture. These models integrate gauge equivariance into both convolutional layers and bilinear operations used to compute Wilson loops, which are gauge-invariant observables from lattice gauge theory. This framework was subsequently extended in [19] to include global lattice symmetries, such as rotations and reflections, thereby achieving a comprehensive symmetry-aware neural network design.